Some summation formula for multibasic hypergeometric series

Udai Bahadur Singh, Lucknow, India

Keywords: Multibasic hypergeometric series, transformation formulae and partial summation formulae AMS Subject classification: 33D65, 33D20

Abstract: In this paper, we derive a multibasic partial summation formulae and deduce, from it, a number of interesting special cases including certain results of Verma, Gasper and also a q-analogue of the finite form of Ramachandra’s result.